Circular No. 6468

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SMC X-1 S. N. Zhang, C. R. Robinson, R. B. Wilson, B. A. Harmon, and D. M.

Scott, on behalf of the BATSE/CGRO team; and R. Remillard, on behalf of the ASM/RXTE team, report: “Quasiperiodic flux variations have been observed from the 0.7-s pulsar SMC X-1 with a period around 60 days, in the energy bands 1.3–12 keV (ASM/RXTE) and 20–100 keV (BATSE/CGRO), confirming the HEAO-1 results in 1977–1978 (Gruber and Rothschild 1984, Ap.J. 283, 546). A sequence of five outbursts between early Oct. 1995 and mid-Aug. 1996 is apparent in the BATSE/CGRO lightcurve, of which the last three coincide with those observed from ASM/RXTE (the first two of the three ASM outbursts have been reported by Levine et al. 1996, Ap.J. 469, L33). The peak flux is ∼ 30 mCrab in both bands. No significant outbursts were observed in the BATSE data between Apr. 1991 and Oct. 1995, though a weak level of emission at ∼ 10–20 mCrab was sometimes detected. Eclipses with a duration of ∼ 20 percent of the orbital period of 3.9 days are observed in both data sets. The current outburst was detected ∼ Aug. 27.”

NOVA CRUCIS 1996 Precise position from G. J. Garradd, Loomberah, N.S.W., obtained on

Sept. 3.409 UT (0.25-m f/4.1 reflector + CCD): α = 12^{h}10^{m}31^{s}.40, δ = −61^{o}45′09′′.6 (equinox 2000.0), mag R = 9.9.

V635 CASSIOPEIAE G. M. Hurst, The Astronomer, reports that G. Poyner (Birmingham,

U.K.) observed a short-lived visual outburst of V635 Cas (= 4U 0115+634) occurring some days after the detection of the reported hard-x-ray flux on Aug. 10 (IAUC 6450), as indicated by the following magnitude estimates: Aug. 8.008 UT, 15.0; 11.956, 15.1; 17.027, 14.7; 18.991, 14.3; 21.979, 14.3; 25.897, 14.6.

COMET C/1995 O1 (HALE-BOPP) K. Jockers, T. Credner, N. Karpov, and A. Sergeev report that, using

the 2-m telescope (+ 2-channel focal reducer of the Max Planck Institute for Aeronomy, Katlenburg-Lindau) of the Centre AMEI, Kiev, at Mount Terskol, they observed the ejection of a CO^{+} cloud into p.a. 130^{o} on Aug. 17.8 UT. On the following nights, the CO^{+} tail gradually subsided.

1996 September 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green